OpenStudy (anonymous):

Please help with this question? im confused! You have 15 red checkers and 5 black checkers. Checkers are selected one at a time with replacement. Each time you record the checker color that is selected. Find the probability of selecting a red checker exactly 4 times in 10 selections.

OpenStudy (anonymous):

binomial probability for this one

OpenStudy (anonymous):

the probability of selecting a red checker on any one pick is \(\frac{15}{20}=\frac{3}{4}\) and the probability of selecting a black one is \(\frac{1}{4}\) you want exactly 4 red in 10 selections

OpenStudy (anonymous):

see I don't understand that. That's why im confused on it. Would it start off with 10c^7?

OpenStudy (anonymous):

i can explain but it will take a second

OpenStudy (anonymous):

i assume you mean \(_{10}C_4\) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

first off, do you get the \(\frac{3}{4}\) part, and the \(\frac{1}{4}\) part?

OpenStudy (anonymous):

Is that the 4 times you could select a red checker outta the 10?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that is the probability you get a red checker

OpenStudy (anonymous):

the ratio of the red checkers to the total number of checkers is \(\frac{15}{20}=\frac{3}{4}\) so the probability you get a red checker on any ONE pick is \(\frac{3}{4}\) and the probability you get a black one is \(\frac{1}{4}\)

OpenStudy (anonymous):

so the beginning of it would look like 10C4 * (3/4)^7 * (1/4)^3

OpenStudy (anonymous):

except that you want 4 red and 6 black

OpenStudy (anonymous):

so it would be \[\large _{10}C_4\left(\frac{3}{4}\right)^{\color{red}4}\left(\frac{1}{4}\right)^6\]